OpenStudy (anonymous):

What is the simplest form of the quotient? ^4√810 over ^4√2

OpenStudy (anonymous):

step one is to divide by 2 and get \[\sqrt[4]{405}\]

OpenStudy (anonymous):

ok i did that now...

OpenStudy (anonymous):

factor 405 may not be obvious, but by some miracle \(405=3^4\times 5\)

OpenStudy (anonymous):

this means \[\sqrt[4]{405}=\sqrt[4]{3^3}\sqrt[4]{5}=3\sqrt[4]{5}\]

OpenStudy (anonymous):

ok so then my answer is \[3 ^4\sqrt{5}\]